### v4

# Data-Copilot: Bridging Billions of Data and Humans with Autonomous Workflow

Abstract
Various industries such as finance, meteorology, and energy produce vast amounts of heterogeneous data every day. There is a natural demand for humans to manage, process, and display data efficiently. However, it necessitates labor-intensive efforts and a high level of expertise for these data-related tasks. Considering large language models (LLMs) showcase promising capabilities in semantic understanding and reasoning, we advocate that the deployment of LLMs could autonomously manage and process massive amounts of data while interacting and displaying in a human-friendly manner. Based on this, we propose Data-Copilot, an LLM-based system that connects numerous data sources on one end and caters to diverse human demands on the other end. Acting as an experienced expert, Data-Copilot autonomously transforms raw data into multi-form output that best matches the user’s intent. Specifically, it first designs multiple universal interfaces to satisfy diverse data-related requests, like querying, analysis, prediction, and visualization. In real-time response, it automatically deploys a concise workflow by invoking corresponding interfaces. The whole processes are fully controlled by Data-Copilot, without human assistance. We release Data-Copilot-1.0 using massive Chinese financial data, e.g., stocks, funds and news. Experiments indicate it achieves reliable performance with lower token consumption, showing promising application prospects.
Figure 1: Data-Copilot is an LLM-based system for data-related tasks, bridging billions of data and diverse user requests. It designs versatile interface tools for data management, invocation, processing, and visualization. Upon receiving a complex request, Data-Copilot autonomously invokes self-design interfaces and constructs a workflow to fulfill human intent. Lastly, raw data is converted into graphics, tables, and text forms tailored to human tastes.
1 Introduction
In the real world, vast amounts of heterogeneous data are generated every day across various industries, including finance, meteorology, and energy, among others. This wide-ranging, multiform data encapsulates critical insights that could be leveraged for a host of applications, from predicting financial trends to monitoring energy consumption.
Recently, the advancement of large language models (LLMs) (Chowdhery et al., 2022; Zhang et al., 2022; Zeng et al., 2023; Touvron et al., 2023), particularly the ChatGPT (OpenAI, 2022) and GPT-4 (OpenAI, 2023a), has paved the way for advanced AI systems. Leveraging chain-of-thought (Wei et al., 2022c; Kojima et al., 2022; Gao et al., 2022b; Wang et al., 2023c), reinforcement learning from human feedback (Ouyang et al., 2022; Wang et al., 2022b), and instruction-following (Wei et al., 2022a; Iyer et al., 2022; Chung et al., 2022), LLMs demonstrate remarkable abilities in dialogue, reasoning and planning. However, in the face of the sheer magnitude and complexity of data, LLMs are confronted with the colossal challenge of managing, processing, and displaying data.
To achieve this, several challenges must be addressed: (1) From a data perspective: employing LLMs for directly reading and processing massive data is not only impractical but also poses the potential risks of data leakage. (2) From the model perspective: due to limitations of the input context, LLMs are unable to process large-scale data at once and are also not adept at handling numerical computations. (3) From the task perspective: many data-related tasks are intricate and require a combination of many operations, like retrieval, computations, and table manipulations, and present their results in multiple formats, e.g., images, tables, and text, which exceeds the capabilities of LLMs. These challenges significantly constrain the application of LLMs in data-related tasks.
Recently, many works have embarked on valuable explorations. LiDA (Dibia, 2023) and GPT4-Analyst (Cheng et al., 2023) focus on automated data exploration. Sheet-Copilot (Li et al., 2023b), BIRD (Li et al., 2023c), DIN-SQL (Pourreza & Rafiei, 2023) and DB-GPT (Xue et al., 2023) apply LLMs to Text2SQL. However, our focus is more on how to better process and present data based on human preference, rather than merely retrieving it from databases. Additionally, while SQL is convenient, it can not directly satisfy common data analysis needs such as prediction and visualization.
In light of this, we contemplate how to integrate LLMs into every facet of data-related tasks, not just as a Text2SQL solver. More specifically, we explore how LLMs can efficiently and autonomously process vast amounts of data to meet diverse human needs, without the necessity of direct data access. To pursue the answer, we trace back to the origins of data science in the 1970s. The pioneer of data science, Peter Naur (Turing Award winner in 2005), defined data science as follows:
Data science is the science of dealing with data and processing large amounts of data. Humans as sources and receivers of data. The data must be chosen with due regard to the transformation to be achieved and the data processing tools available. —Naur, 1974
His statement elucidates how humans accomplish data-related tasks: humans create various tools that are easy to use and then transform the data into a human-friendly format. Similarly, LLMs also should play a more intelligent role like humans: designing tools for data management, and invoking tools for processing and presenting data.
Additionally, numerous LLM-based agent methods have recently emerged (Wu et al., 2023a; Huang et al., 2023; Chen et al., 2023b; Hong et al., 2023; Wu et al., 2023b). These LLM-based agents showcase the potential of LLMs in meeting human needs through code as an intermediary.
Inspired by them, we advocate utilizing the coding capabilities of LLMs to create diverse and powerful tools, thereby facilitating the management and utilization of vast amounts of data. Based on this, we propose Data-Copilot, a reliable data agent that manages vast data and dynamically fulfills a wide spectrum of human needs. As shown in Figure 1, Data-Copilot autonomously designs versatile interface tools for data management, and dispatch these interface tools step by step for diverse human requests, e.g., data analyzing, forecasting, and visualization, forming a data-to-human workflow.
As an LLM-based agent, Data-Copilot is versatile. It seamlessly connects an array of data sources from various domains on one end, while adeptly catering to diverse user demands on the other. Furthermore, Data-Copilot is scalable. It can continuously extend its capability boundaries by developing new interface tools for emerging new data and requests. Our system comprises two processes:
  • Interface Design: Data-Copilot adopts an iterative self-request process to fully explore the data, mimicking a broader range of human need scenarios. As shown in Figure 1, starting from a few seed requests, it self-generates a large number of requests and then abstracts them into generalized interface tools for efficient utilization. These interfaces empower Data-Copilot to efficiently perform data querying, processing, forecasting, and visualization.
  • Interface Dispatch: When a user request is received, Data-Copilot first parses the user intention and then plans an interface invoking workflow for the user request. This step-by-step invocation can efficiently fulfill user requirements while minimizing token overhead.
Figure 2: Overview of Data-Copilot. Interface Design: We devise a self-request process allowing LLM to generate sufficient requests from a few seed requests autonomously. Then LLM iteratively designs and optimizes interfaces based on generated requests. These interfaces are described in natural language, making them easily scalable and transferable across different platforms. Interface Dispatch: Upon receiving user requests, LLM plans and invokes interface tools based on self-design interface descriptions and in-context demonstrations. This allows for the deployment of a logical workflow that fulfills user demands and presents the results to the user in multi-form.
Incorporating two phases, Data-Copilot achieves autonomous data management and utilization without the need for direct data access. Our contributions can be summarized as follows:
  1. 1.
    To efficiently handle data-related tasks, we design a universal system, Data-Copilot, that connects massive data sources to diverse user tastes. It leverages LLM for more efficient data management and utilization, reducing tedious labor.
  2. 2.
    Data-Copilot can autonomously manage, process, analyze, and visualize data. For a human request, it transforms raw data into informative results that best match the user’s intent.
  3. 3.
    As a reliable assistant, Data-Copilot abstracts a multitude of common needs into various practical interface tools. When faced with real-time requests, it invokes these related interfaces sequentially or in parallel for problem-solving.
  4. 4.
    We evaluate Data-Copilot on massive Chinese financial data, e.g., stocks, funds, macroeconomics, and news. The excellent success rate and lower computational cost showcase its potential and application prospects.
2 Data-Copilot
Data-Copilot is an LLM-based agent that accomplishes automated management, invocation, and processing of a large number of data by self-design sophisticated interface tools. Besides, Data-Copilot adeptly comprehends user intentions, and invokes well-designed tool interfaces for request solving, significantly reducing the need for tedious labor.
As shown in  Figure 2, the LLM plays the role of a designer (Section 2.1) to independently design various tool interfaces to for data management and utilization. Besides, LLM also acts as a dispatcher (Section 2.2), invoking generated interfaces and deploying the workflow for problem-solving. Through the integration of two stages, Data-Copilot manages to efficiently handle a large volume of data-related tasks and caters to various user needs.
2.1 Interface Design
Just as Peter Naur pointed out, designing a set of versatile interface tools that are easy to manage and use is critical. Therefore, Data-Copilot designs a plethora of interfaces as tools, where each interface is composed of natural language (functional description) and code (implementation), responsible for data acquisition, processing, and visualization.
As shown in Figure 2, beginning with some seed requests, LLM autonomously mimics and generates a large number of requests. Then LLM designs common interfaces for these requests, representing a set of universal needs. By gradually refining and expanding their functionalities, these interfaces progressively cover a more diverse array of needs. Ultimately, the LLM translates these interfaces into executable and callable code modules.
Explore Data by Self-Request
The design of the interface depends on two aspects: what kind of data is available and also what kind of demand the user proposes. Inspired by (Wang et al., 2022a; Dibia, 2023), Data-Copilot first autonomously explores the data to mine more requests. Then it proceeds to design the interfaces for these requests.
Specifically, we generate a parsing file for each data source to help LLM understand the data. Each file includes a description of the data, the access method, the data schema (the name of each column), and a usage example. This process does not require too much labor, as data providers usually offer a description of the data and access methods.
Then we adopt a self-request process to explore data, i.e. the parsing files and a few seed requests are fed into the LLM and then LLM is instructed to generate more diverse requests by mimicking. This process aims to explore an extensive array of data sources while covering diverse user requests. In Figure 2, LLM generates four new requests based on the two seed requests and data parsing files. The detailed prompts are shown in Section F.1.
Interface Definition
In this step, Data-Copilot creates various interface tools to fulfill the previously generated requests. Specifically, we feed all data parsing files and all interfaces stored in the interface library (empty at first iteration) into LLM as a prompt. LLM needs to resolve each request one by one. When addressing each request, LLM can use existing interfaces from the interface library or redefine a new interface to fulfill the task.
Concretely, if existing interfaces are insufficient to address the current request, LLM can create a new interface: Interface1={Name:"", Function description:"", Input and Output:""}... After that, LLM proposes a feasible solution using these self-design or existing interfaces: Solution="These three interface functions will meet your needs. You should first use the interface: {getBankData} to acquire data, then...". The prompts and cases are shown in Sections F.1 and G.1.
Importantly, Data-Copilot only uses natural language to define the interface’s functions and their arguments, without considering the specific implementation, which is grammar-free. This process is similar to software architecture design by a human programmer. It allows Data-Copilot to focus on designing the layout and relation of the interfaces with different functionalities, rather than complying with programming grammar.
Interface Merging
To make the interface more universal, Data-Copilot considers whether each newly designed interface can be merged with the existing ones in the library. Specifically, when a new interface is designed, Data-Copilot also checks whether this interface is similar to the previous ones, in terms of functionality, arguments, etc. Two similar interfaces are merged to create a new and generalized interface. As shown in Figure 2, two interfaces {Interface: Query-GDP} and {Interface: Query-CPI} are merged into one {Interface: Query-GDP-CPI}. This merging process is similar to software developers wrapping similar modules into a new one.
After this process, a large number of similar interfaces are merged. Each interface within the library possesses distinct functionalities, which are adeptly suited for real-time response applications.
Interface Implementation
Data-Copilot iteratively processes each request through interface definition and interface merging. Eventually, when all the requests can be satisfied by these self-design interfaces, Data-Copilot transforms these interface tools into executable and callable code modules.
The whole interface design process is offline. As shown in Figure G3, Data-Copilot automatically produces five types of interfaces in the interface library: data acquisition, processing, prediction, visualization, and DataFrame manipulation. It transforms repetitive and tedious labor into an automated process, and also can effortlessly add new interfaces for additional requests or new data sources. Besides, Data-Copilot easily switches to other programming platforms and databases by simply re-generating the implementation code for the interface, demonstrating excellent scalability.
2.2 Interface Dispatch
In the first stage, Data-Copilot designs a variety of generic interface tools for data acquisition, processing, and visualization. These interfaces are versatile and easy to use, improving the efficiency of real-world response. These interfaces are invoked in this phase and form an efficient workflow for real-time requests (Figure 2). A detailed prompt and cases are shown in Sections F.2 and G.2
Intent Analysis
To accurately comprehend user requests, Data-Copilot first parses the time, location, data object, and output format of user requests, which are critical to data-related tasks. For example, if the question is: "I want to compare the GDP and CPI trend in our area over the past five years", Data-Copilot parses it as: "Draw a line chart of China’s national GDP and CPI per quarter from May 2017 to May 2023 for comparison". To achieve this, we first invoke an external API to obtain the local time and network IP address, then feed this external information into LLM along with the original request to generate the parsed result.
Planing Workflow
Once the user’s intent is accurately understood, Data-Copilot plans a reasonable workflow to process the user’s request. We feed the user request and related interface descriptions into LLMs for the automatic deployment of workflows. Which interfaces should be invoked, and in what sequence, is entirely determined by Data-Copilot, based on the user’s request and the interface.
To achieve this, Data-Copilot generates a response in JSON format representing each step of the scheduling: ###Interface Call: {step1={"arg":"", "function":"", "output":"" ,"description":""}, step2={..}, ..}. Please refer to Section F.2 for a detailed prompt. Data-Copilot meticulously orchestrates the scheduling of the interface in either a sequential or parallel manner.
For instance, the first request in Figure 2: "Compare the return of CSI 300, GEM Index and CSI 1000 Index this year", Data-Copilot first plans five steps in its workflow. In the first two steps, it dispatches {Data Acquisition}, {Calcuate Return} in parallel for three indices. The last three steps successively invoke the {Visualization} to sequentially plot three indices on the same canvas. In the second case, Data-Copilot deploys a loop workflow to sequentially calculate the indicators of 50 constituent stocks by calling the {Loop and Rank}.
Multi-form Output
Upon the deployment and execution of the workflow, Data-Copilot yields the desired results in the form of graphics, tables, and descriptive text. Additionally, it also provides a comprehensive summary of the entire workflow. This systematic summary not only provides clear and concise results but also sheds light on the steps taken to achieve those results, thereby enhancing transparency and understanding of user requests.
As the example shown in Figure G2, the request is "Forecasting China’s GDP growth rate...". Data-Copilot first interprets the user’s intent based on local time. Then, it deploys a three-step workflow: 1) Invoking {get-GDP-data} interface to acquire historical GDP data. 2) Invoking {predict-next-value} interface for forecasting. 3) Visualizing the output.
3 Qualitative Evaluation
3.1 Environment and Data Sources
We evaluate Data-Copilot utilizing Chinese financial market data, encompassing massive stocks, funds, economic data, real-time news, and company financial data. To closely mimic real-world scenes, we adopt the data interfaces (APIs) provided by Tushare [^1] https://tushare.pro/ to access data, which are commonly used by many financial companies. The vast and diverse array of financial time-series data enables a comprehensive evaluation of Data-Copilot’s performance. In the first phase, we use the gpt-4-0613222https://platform.openai.com/ for interface design and gpt-3.5-turbo-0613 for interface dispatch.
3.2 Seeds and Datasets
Human-proposed Request as Seed We invite 10 students in economics to submit 50 requests each. These requests should satisfy two requirements: 1) cover a wide range of common needs, including stocks, funds, and corporate. 2) involves different levels of difficulty. Then we filter out highly similar requests by computing the sentence embedding of each request. Lastly, we retain 173 high-quality requests, which are used as the seeds for Interface Design phase of the Data-Copilot.
Self-Request and Testing Benchmark As mentioned in Section 2.1, we employed a self-request strategy to expand the request set. Specifically, we feed 173 manually proposed seed requests along with all data descriptions and schema into GPT-4. We instruct GPT-4 to generate massive requests, covering diverse data. In this phase, we generate a total of 6480 requests. Then we compute the semantic similarity between all requests and filter out those with high similarity. Lastly, we retain 3547 requests. We randomly select 3000 requests from the total set for subsequent interface design. Meanwhile, the remaining 547 requests are treated as a test set. We manually annotate the final answers, task types, and workflow types for each testing instance. The statistics of three dataset are shown in Tables 1 and C1.
3.3 Qualitative Results
The whole system is shown in Figure G9, where Data-Copilot dynamically displays the intermediate process (Solving Step) and the final results (bar, chart, text) for the user request. The invocation workflow usually consists of multiple steps, each of which invokes one or multiple interfaces. For example, Data-Copilot addresses the request in Figure 2 ("Compare the return of ...") via a five-step procedure, wherein the first two steps invoke the same interface in parallel.
Data-Copilot designs many versatile interface tools through a self-request. As shown in Figure G3, each interface is clearly defined and versatile, including data acquisition, processing, dataframe manipulation, and visualization. These diverse interfaces endow the LLM with the capability to address complex requests. For instance, get-stock-prices-data interface integrates the functions of acquiring data in daily, weekly, and monthly frequencies, controlled by argument: freq, which is very similar to the process of programmer development.
Data-Copilot deploys workflows with different structures. Data-Copilot invokes the most appropriate interface and forms an efficient workflow with different structures, e.g., parallel, serial, or loop workflows, tailored to user requests. As shown in Figures G4 and G5, it employs a loop workflow for the requests: "The year-on-year net profit growth rate of the SSE50 ..." by invoking Loop-Rank interface. It iteratively calculates the financial indicators for each stock. Please refer to Appendix E for more visualization cases.
4 Quantitative Evaluation
We quantitatively assess Data-Copilot in processing real-time user requests, where it could invoke the interfaces designed in the first phase.
Table 1: Left: Statistics on human-proposed requests and Self-Request dataset. Right: Detail Statistics on different request types and workflow structures.
Source #Cases Form Usage
Human-proposed 173 Query Seed
Self-Request 3000 Query Interface Design
Self-Request + Human-annotated 547 Query, Ans Testing
Request Types Workflow Types
Sequential Parallel Complex Overall
Stock 79 106 87 272
Fund 55 36 22 113
Corp. 97 30 4 131
Other 6 12 13 31
Total 237 184 126 547
4.1 Experiments Settings
Baselines 1. Global Planning (GP) (Shen et al., 2023; Qin et al., 2023a): We instruct the LLM to directly generate solving codes for user requests. 2. ReAct (Yao et al., 2022): The LLM resolves the user request using neural language and code step-by-step. 3. Reflexion (Shinn et al., 2023): LLM can also reflect on the feedback from the compiler at each step, progressively fulfilling user requests. For the Reflexion strategy, we limit it to a maximum of two rounds of reflection. 4. Multi-agent collaboration (Wu et al., 2023b; Hong et al., 2023; Liang et al., 2023a): We adopt three LLM-based agents, i.e., two coders and a manager, handling user requests simultaneously and outputting the final solution after mutual discussion. All methods are evaluated on the 547 testing instances (Section 3.2) three times. A detailed prompts of baselines is shown in Section F.3.
Metrics We design three metrics for solution code and answer: Exec@1 Pass@1 and Efficiency. Exec@1 evaluates whether the code can be executed by a compiler without errors. Pass@1 measures the correctness of the final answer. Considering the diverse output formats, e.g., images, text, numbers, and tables, we employ GPT-4 to automatically assess the answer. If the output is an image, we use gpt-4-vision-preview to compare the generated image to the label. Lastly, we calculate the total number of tokens consumed by the input prompt and output response, which denotes the solving efficiency of each method.
Table 2: All methods use gpt-3.5 as based LLM.
Methods Exec@1(%) Pass@1(%) #Token
GP 50.8 28.5 823.6
ReAct 63.4 44.1 1515.9
Reflexion 75.7 59.0 2463.5
Multi-Agent 73.6 57.4 2835.2
Data-Copilot 79.9 70.2 561.2
Table 3: All samples are divided into three subsets based on its complexity. We compute Pass@1.
Methods Sequential Parallel Complex Overall
GP 42.4% 29.0% 1.6% 28.5%
ReAct 56.3% 48.6% 14.6% 44.1%
Reflexion 67.8% 55.1% 48.1% 59.0%
Multi-agent 63.2% 64.7% 35.5% 57.4%
Data-Copilot 71.8% 70.0% 67.1% 70.2%
4.2 Comparison Results
As shown in Table 3, Data-Copilot surpasses all baselines in success rate (Pass@1), Exec@1, and Efficiency, achieving 11.2% improvements.
Data-Copilot significantly reduces the risk of failure. Compared to the best baseline (Reflexion), our method increases the success rate from 59% to 70.2%. In most cases, our system merely needs to invoke pre-designed interface tools and generate corresponding arguments to address user requests, significantly reducing the risk of errors. In contrast, baseline strategies must construct all code from scratch, which makes it easy to overlook certain details or make mistakes. Besides, the difference between Pass@1 and Exec@1 also indicates that baseline strategies often miss certain steps, whereas Data-Copilot can fully interpret user requests and methodically complete tasks step by step.
Data-Copilot reduces repetitive overhead for common requests. The consumed tokens represent computational overhead. Data-Copilot only consumes 561 tokens, which is just 30% of the consumption compared to baselines. Specifically, the output of Data-Copilot is very concise, only containing interface names and arguments. The primary token consumption lies in the input prompt which contains related tool descriptions. In contrast, baseline strategies may consume a substantial amount of tokens since they need to output a complete solving code for each request. In the real world, numerous requests are similar or even repetitive. Data-Copilot can reduce repetitive overhead by abstracting common demands into interface tools. However, baseline methods are compelled to repetitively generate code for each request.
4.3 Analysis
Data-Copilot exhibits a pronounced advantage on more challenging tasks. To better analyze the problem-solving capabilities of Data-Copilot, we categorize our dataset into three subsets based on the complexity of the workflow: Sequential, Parallel, and Complex. As shown in Table 3, for sequential tasks, the baselines and our method perform similarly, with only a 4% improvement. However, for the more challenging parallel or complex tasks, our improvements are substantial, with enhancements of 5.3% and 19%, respectively. We find that baseline methods often struggle to fully implement complex functionalities and may omit some critical steps, such as loop processing. In contrast, our methods effectively handle these tasks by invoking versatile interfaces and deploying workflow. These findings elucidate the advantages of Data-Copilot, which reduces task complexity.
We can deploy Data-Copilot locally using small LLMs with commendable performance. We analyze the performance of Data-Copilot with different LLMs. As shown in Figure 4, we test GPT family APIs and various open-source models. We observe GPT-4 surpasses all LLMs, achieving 75.6% Pass@1, while many open-source LLMs also achieve commendable results. For instance, codellama-13B achieves a 50.8% Pass@1, almost similar to the best baseline in Table 3 (Reflexion with gpt-3.5-turbo). This indicates that Data-Copilot can be deployed locally at a relatively low cost while maintaining commendable performance.
4.4 Human Evaluation
To assess whether the solutions align with human preferences, we perform human evaluations across two dimensions: rationality and conciseness. Specifically, we collect the solutions of Data-Copilot and Reflexion. Then we invite experts to compare which solution is better in terms of rationality and conciseness (detail in Appendix C). We present results in Figure 4. It demonstrates that our solution significantly outperforms the competitor in terms of rationality. Specifically, Data-Copilot prevails in 47.4% of instances, trailing behind the competitor in only 23.5%. Besides, as for conciseness, Data-Copilot lags slightly behind its competitor, with 27.8% wins against 30% losses. This is attributed to the fact that these pre-designed interfaces invoked by Data-Copilot are quite universal, accommodating a wide array of functions, which makes our system less concise. The results of human evaluation are consistent with the previous assessment. Although Data-Copilot sacrifices a bit of succinctness, it can generate more rational and correct solutions.
Figure 3: Results of Data-Copilot with various LLMs.
LLM Pass@1 (%) #Token
gpt-4 75.6 561.2
gpt-3.5-turbo 70.2 559.3
Llama-2-70B-chat 49.3 623.6
codellama-13B 50.8 507.4
text-davinci-003 47.5 495.9
Llama-2-13B-chat 36.9 443.8
Vicuna-13B-v1.5 33.2 502.8
Figure 4: Win rate for Data-Copilot Vs. Reflexion on human preferences.
5 Conclusion
We propose a universal framework, Data-Copilot, to address a wide array of data-related tasks. It acts as a bridge between numerous heterogeneous data and humans, effectively managing, processing, and presenting data according to human tastes, reducing the dependence on tedious labor and expert knowledge. Acting like an experienced expert, Data-Copilot autonomously designs universal interface tools for various types of data and potential user demands and invokes these interfaces in real-time response for more reliable problem-solving.
6 Acknowledgement
This work is supported by the National Natural Science Foundation of China (No. 62376245), the Key Research and Development Program of Zhejiang Province, China (No. 2024C01034), the project of the Donghai Laboratory (Grant no. DH-2022ZY0013), and MOE Engineering Research Center of Digital Library.
Appendix
Appendix A Discussion
A.1 What is Data-Copilot?
  • Data-Copilot is an efficient LLM-based data management agent, that processes massive data and rapidly responds to real-time human requests.
  • It contains interface design and invocation phases. Interface design involves a self-exploration phase for data sources to create many versatile interfaces.
  • For interface invocation, it parses human real-time requests and invokes the designed interfaces to retrieve and process relevant data and generates multi-form outputs, e.g., text, charts, and tables.
A.2 Why Data-Copilot is designed in two stages?
  • Reduces redundant generation: People often have a large number of similar or repetitive query needs, e.g., many people are interested in seeing which stock has the largest gain today. For these repetitive query requests, employing the LLM to write code from scratch for each user request would be highly inefficient and a significant waste of computational resources. To avoid inefficient and repetitive code generation for numerous similar requests, we abstract common demands into generic interfaces through data self-exploration in advance (Phase-1). For real-time requests (Phase-2), it only needs to invoke these interfaces, reducing redundant generation.
  • Performs better and consumes fewer tokens: Compared to a complete code, it only needs to invoke the relevant interfaces, which is simpler and less prone to errors. Our results show that it performs better and consumes fewer tokens.
  • Stronger interpretability: Compared to complex code, interface workflow is easier for humans to read and inspect.
Appendix B Experiments Details
During the interface design phase, we instruct LLM to generate Python code for each interface in the interface implementation. We filter out interfaces that cannot run properly at each iteration. For the interface Dispatch phase, we carry out hierarchical planning to improve efficiency: upon receiving a request, Data-Copilot first determines the type of data task involved (stock task, fund task, etc.) and then loads the corresponding interface description, followed by interface planning using the corresponding interface. We set the temperature to 0.2 for more consistent results.
Appendix C Human Evaluation Details
We collect the solutions of Data-Copilot and Reflexion on 173 human-proposed requests. Subsequently, we involve experts to evaluate which of the two solving codes is more rational and concise. The criteria for expert are as follows:
Rationality:
1) Accurately identifies data of interest to the user.
2) Processes the data logically.
3) Output data in a format that meets human needs, with aesthetic and concise presentation.
Conciseness:
1) Only extracts and computes data relevant to the users interest.
2) Code is clear, concise, and well-structured.
3) The solution process is devoid of repetition and redundancy, ensuring a straightforward and intuitive approach.
Experts are required to evaluate which of the two solutions is better based on these principles, and then we calculate the win rate of Data-Copilot against the Reflexion.
Figure C1: We count the keywords for each request type (Stock, Corp., Fund, Others) in the test set.
Appendix D Related Works
In the recent past, breakthroughs in large language models (LLMs) such as GPT-3, GPT-4, PaLM, and LLaMa (Brown et al., 2020; Chowdhery et al., 2022; Zhang et al., 2022; Zeng et al., 2023; Touvron et al., 2023; Ouyang et al., 2022; OpenAI, 2023b; Wei et al., 2022b; Wu et al., 2023c) have revolutionized the field of natural language processing (NLP). These models have showcased remarkable competencies in handling zero-shot and few-shot tasks along with complex tasks like mathematical and commonsense reasoning. The impressive capabilities of these LLMs can be attributed to their extensive training corpus, intensive computation, and alignment mechanism (Ouyang et al., 2022; Wang et al., 2022b; a).
LLM-based Agent Recent studies have begun to explore the synergy between external tools and large language models (LLMs). Tool-enhanced studies  (Schick et al., 2023; Gao et al., 2022a; Qin et al., 2023a; Hao et al., 2023; Qin et al., 2023b) integrate external tools into LLM, thus augmenting the capability of LLMs to employ external tools. Several researchers have extended the scope of LLMs to include the other modality (Wu et al., 2023a; Surís et al., 2023; Shen et al., 2023; Liang et al., 2023b; Huang et al., 2023). In addition, there are many LLM-based agent applications (Xie et al., 2023), such as CAMEL (Li et al., 2023a), AutoGPT [^3] https://github.com/Significant-Gravitas/Auto-GPT, AgentGPT [^4] https://github.com/reworkd/AgentGPT, BabyAGI [^5] https://github.com/yoheinakajima/babyagi, BMTools [^6] https://github.com/OpenBMB/BMTools, LangChain [^7] https://github.com/hwchase17/langchain, Agentverse (Chen et al., 2023b), Autoagent (Chen et al., 2023a), MetaGPT (Hong et al., 2023), AutoGEN (Wu et al., 2023b), etc. Most of them are focused on daily tools or code generation and do not consider the specificity of data-related tasks. Except for learning to operate the tools, several contemporaneous studies (Cai et al., 2023; Qian et al., 2023) have proposed to empower LLMs to create new tools for specific scenarios like mathematical solving and reasoning. These impressive studies have revealed the great potential of LLM to handle specialized domain tasks.
Applying LLM To Data Science Apart from these studies, the application of large models in the field of data science has garnered significant interest among researchers. FLAME (Joshi et al., 2023) investigates the feasibility of using NLP methods to manipulate Excel sheets. StructGPT (Jiang et al., 2023) explore reasoning abilities of LLM over structured data. LiDA (Dibia, 2023) and GPT4-Analyst (Cheng et al., 2023) focus on automated data exploration. Besides, many reseraches (Liu et al., 2023; Chang & Fosler-Lussier, 2023; Dong et al., 2023), like Sheet-Copilot (Li et al., 2023b), BIRD (Li et al., 2023c), DAIL-SQL (Gao et al., 2023), DIN-SQL (Pourreza & Rafiei, 2023), DB-Copilot (Wang et al., 2023b), MAC-SQL (Wang et al., 2023a), ChatDB (Hu et al., 2023), SQL-PaLM (Sun et al., 2023), and DB-GPT (Xue et al., 2023) apply LLMs to Text2SQL and table rasoning. Chain-of-Table (Wang et al., 2024) proposes a step-by-step reasoning strategy based on the table.
Distinct from these approaches, our Data-Copilot system is different in the following aspects:
(1) Different Task Settings: Unlike Text2SQL methods, Data-Copilot focuses more on how to utilize data and present it according to user preferences, rather than retrieving data from databases. We assume that Data-Copilot has access to a vast amount of data, which is available via API.
(2) Richer Usage Scenarios: Data-Copilot needs to meet a diverse range of human needs, such as forecasting and plotting. Therefore, it employs programmable language like Python for data processing, offering more varied use cases than SQL.
(3) Distinct Implementation Approach: Data-Copilot designs its tools for data processing, rather than relying on SQL or pre-existing tools. This approach is more flexible and can meet complex user requirements.
Appendix E Visualization
We provide several cases in this section to visualize workflow deployed by Data-Copilot, which includes queries about diverse sources (stocks, company finance, funds, etc.) using different structures (parallel, serial, and loop Structure).
Different Structures
As shown in Figure G4,G5, Data-Copilot deploys different structural workflows based on user requirements. In Figure G4, the user proposes a complex request, and Data-Copilot deploys a loop structure to implement the financial data query of each stock, and finally outputs a graph and table in parallel. In Figure G5, Data-Copilot proposes a parallel structure workflow to meet user request (the demand for comparison in user request) and finally draws two stock indicators on the same canvas. These concise workflows can cope with such complex requests well, which suggests that the interface design and dispatch process of Data-Copilot are rational and effective.
Diverse Sources
Figure G6, G7, G8 demonstrate that Data-Copilot is capable of handling a large number of data sources, including stocks, funds, news, financial data, etc. Although the formats and access methods of these data types are quite different, our system efficiently manages and displays the data through its self-designed versatile interface, requiring minimal human intervention.
Appendix F Detailed Prompts
We provide detailed prompts for our Data-Copilot and baselines. Specifically, for the interface design, we provide detailed prompts in Section F.1 for four phases: Self-Request, Interface Definition, Interface Merging, and Interface Implementation. For the interface dispatch, we also provide prompts for three key procedures in Section F.2: Intent Analysis, Task Selection, and Planning Workflow. Additionally, we outline prompts for all baselines in Section F.3: Global Planning, ReAct, Reflection, and Multi-Agent Collaboration strategies.
F.1 Prompts For Interface Design
Explore Data by Self-Request Phase
###Instruction: Given some data and its description, please mimic these seed requests and generate more requests. The requests you generate should be as diverse as possible, covering more data types and common needs.
###Seed Request: {request1, request2,...}
###Data files:{
GDP-Data:{
Description: This data records China’s annual and quarterly GDP...,
Access Method: You can access the data by pro.cn-gdp(start-time, end-time, frequency,...), start-time means....,
Output Schema: The data return 9 columns, including quarter: quarter, gdp: cumulative GDP, gdp-yoy: quarterly Year-on-Year growth rate, pi,...,
Usage:{
Example: pro.cn-gdp(start-q=’2018Q1’, end-q=’2019Q3’, frequency=’quarter’),
First Row: {2019Q4 990865.1 6.10, ...},
Last Rows: {2018Q4 900309.5 6.60, ...}
stock_basic:...,
Interface Definition Phase:
###Instruction: You are an experienced program coder. Given a request and some existing interfaces, you should use these interfaces to solve the request or design new interfaces (similar to pseudocode) to resolve my request.
(1) You should define the name, function, inputs, and outputs of the interface. Please describe the functionality of the interface as accurately as possible and do not write any implementation code in the new interface, just design the interface as existing interfaces.
(2) Finally please explain how to resolve my request using your newly designed interfaces or existing interfaces in the neural language.
###Output Format:
Your newly designed interfaces:
Interface1={Interface Name: {name},
Function description: {This interface is to ...},
Input: {argument1: type, argument2: type, ...},
Output: {pd.DataFrame} }
Interface2=....
The solving process for request: {To fullfil this request, I design a interface ...}
###The user request: {input request}
###Data Files: {All data files}
###The existing interfaces: {all interfaces in library}
Interface Merging Phase
###Instruction: Please check that the interface you have designed can be merged with any existing interfaces in the library.
(1) You should merge interfaces with similar functionality and similar input and output formats into a new interface.
(2) You can use parameters to control the different inputs. Finally
(3) Please explain your reason for merging and output all interfaces in the library after merging.
(4) If you dont think a merge is necessary, then just add new interfaces into the existing interface library and output them all.
###New interfaces: {interfaces}
###Existing interfaces: {interface1, interface2, interface3, ..}
###Output Format:
The reasons for merging: {reason}
All interfaces after merging: {interface1, interface2, merged interface3, ...}
Interface Implementation Phase:
###Instruction: Please generate the complete code according to the definition of a given interface and data access methods. The generated code must be in strict accordance with the definition of the interface and the formatting requirements of the arguments.
###Interface: {Input Interface}
###Data Files: {All data files}
F.2 Prompts For Interface Dispatch
Intent Analysis Phase
### Analysis Prompt: Please parse the input request for time, place, object, and output format. You should rewrite the instruction according to todays date. The rewritten new instruction must be semantically consistent and contain a specific time and specific indicators.
### Output Format: Rewritten Request. (Time:%s, Location:%s, Object:%s, Format:%s).
### User Request: Today is {Timestamp}. The user request is {Input Request}.
Please output a Rewritten Request.
Task Selection
###Select Prompt: Please select the most suitable task according to the given Request and generate its task_instruction in the format of task={task_name: task_instruction}. There are four types of optional tasks. [fund_task]: used to extract and process tasks about all public funds. [stock_task]: for extracting and processing tasks about all stock prices, index information, company financials, etc., [economic_task]: for extracting and processing tasks about all Chinese macroeconomic and monetary policies, as well as querying companies and northbound funds, [visualization_task]: for drawing one or more K-line charts, trend charts, or outputting statistical results.
###Output Format: task1={%s: %s}, task2={%s: %s}
###User Request: {Rewritten Request}.
Please output a task plan for this request.
Planning Workflow
###Planning prompt: Please use the given interface (function) to complete the Instruction step by step. At each step you can only choose one or more interfaces from the following interface library without dependencies, and generate the corresponding arguments for the interface, the arguments format should be strictly in accordance with the interface description. The interface in the later steps can use results generated by previous interfaces.
###Output Format:
Please generate as json format for each step:step1={"arg1": [arg1,arg2...], "function1": "%s", "output1": "%s", "description1": "%s"}, step2={"arg1": [arg1,arg2..], "function1": "%s", "output1": "%s", "description1": "%s"}, ending with ###.
###User Request: {Task Instruction}.
Please output an interface invocation for this instruction.
F.3 Prompts For Baselines
Global Planning
###Instruction: You are an artificial intelligence assistant. Given some data access methods and a user request, you should write a complete Python code to fulfill the users request. Your code must completely fulfill all the users requirements without syntax errors!
###User Request: {User request}
###Data files: {All data files}
Please solve the request by Python Code.
Step-by-Step ReAct
###Instruction: You are an artificial intelligence assistant. Given some data access methods and a user request, please think step by step and generate your thoughts and actions for each step, and then finally realize the users request.
###User Request: {User request}
###Data files: {All data files}
###Thought Prompt: Please think about the next action that should be taken to handle the user request.
### {Thought: I need to ...}
###Action Prompt: Based on your previous thoughts, please generate a complete Python code to accomplish what you just planned.
### {Action: def get-data()....}
###Observation Prompt: Please summarize the results of the code execution just now and think about whether this result accomplishes what you planned for this step.
### {Action: Yes, I observed that this function successfully fetched the data...}
Step-by-Step Reflexion
###Instruction: You are an artificial intelligence assistant. Given some data access methods and a user request, please think step by step and then generate your thoughts and actions for each step. After the execution of your current action, you need to reflect on the results until your current plan has been successfully completed. Then you think about the next step and then generate your next action, and finally realize the users request.
###User Request: {User request}
###Data files: {All data files}
###Thought Prompt: Please think about the next action that should be taken to handle the user request.
### {Thought: I need to ...}
###Action Prompt: Based on your previous thoughts, please generate a complete Python code to accomplish what you just planned.
### {Action: def get-data()....}
###Observation Prompt: Please record the compilers return results just now and think about whether this result accomplishes what you planned for this step.
### {Action: No, I observe that the compiler returns an error...}
### Reflection Prompt: Please reflect on the error returned by the compiler and regenerate a new Python code to resolve the issue. If the compilation passes without any errors, reflect on whether the current result is what you planned to do.
### {Action: I revise my solution as follows: def get-data2()....}
Multi-agent collaboration
###Instruction For Manager: You are the manager of the project team and you need to lead your team to fulfill user requests. You have two experienced programmers under you (Programmer-A and -B) and you need to assign them the same or different tasks according to the user request, then organize the discussion, and finally solve the problem.
###Instruction For Agent1/Agent2: You are an experienced programmer. Your team has a colleague who is also a programmer and a manager. You need to write code according to the managers arrangement, discuss it with them, improve your program, and get a consensus conclusion.
###User Request: {User request}
###Data files: {All data files}
-------- Phase1 Discussion for Task Assignment -------
###Task Assignment Prompt For 3 Agents: Now lets start discussing how to solve user problems. The speaking order is manager, A and B. After a round of discussion, the manager decides on a preliminary solution and assigns specific tasks to the two programmers
###Manager Response: {To fulfill the request...}
###Programmer A Response: {Yes, I agree with the manager. We should first,....}
###Programmer B Response: {That’s a good solution. We also should...}
###Manager Response: {After discussion, Programmer A should..., B should ....}
-------- Phase2 Coding and Discussion-------
###Prompts For 3 Agents: Two programmers complete their respective tasks as assigned by their manager. Then the manager organizes a discussion on whether the two options and their running results are reasonable. Lastly, two programmers need to regenerate the code.
###Programmer A Response: {Solution1 Generated by A}
###Programmer B Response: {Solution1 Generated by B}
###Prompts For 3 Agents: Please evaluate two solutions. {Solution1, Result1, Solution2, Result2}
###Programmer A Evaluation: {I think ...}
###Programmer B Evaluation: {I agree with ...}
###Programmer A Regeneration: {Revised Solution1}
###Programmer B Regeneration: {Revised Solution2}
-------- Phase3 Discussion-------
-------- Phase4 Final Decision-------
###Prompts For manager: Please output a complete and correct solving code for user request based on the previous discussions of the group.
###Manager Decision: {Final Solution}
Appendix G Case Study
G.1 Cases for Interface design
---- The First Request: Interface Definition Phase -----
Input Request: Get the daily opening price data of Bank of Hangzhou from Jan 23, 2018 to Mar 13, 2019 and plot the trend.
Response: First, we need to design a function to get the opening price data from the banks database. Then, we need another function to process this data and prepare them for graphical representation. Finally, we need a function to plot the trend chart.
New Design Interfaces:
{Interface1=Interface Name:getBankData, Function:.., Input:.., Output:..}
{Interface2=Interface Name:prepareDataForPlotting, Function:.., Input:.., Output:..}
{Interface3=Interface Name:plotData, Function:.., Input:.., Output:..}
The solving process for request: These three new interface functions will meet your needs. In your program, you would first get the data using getBankData, then process the data using prepareDataForPlotting, and finally use plotData to plot the trend.
---- The Second Request: Interface Definition Phase-------
Input Request: Get the monthly closing price data of Ningde Times from January 23, 2018 to March 13, 2019, and plot the trend.
Response: To fulfill this requirement, we design a new interface function. In addition, I need to use two existing interfaces to fulfill your requirement.
New Designed Interfaces:
{Interface4=Interface Name: getStockData..}
Existing Interfaces
{Interface1=Interface Name:getBankData, ...}
{Interface2=Interface Name:prepareDataForPlotting, ...}
{Interface3=Interface Name:plotData, ...}
The solving process for request: You need to call these functions to meet your needs. First, use new interface {getStockData} to get the data of Ningde Times, then use existing interface {prepareDataForPlotting} to process this data, and finally use {plotData} to plot the trend...
---- Interface Merge Phase -------
The reason for merging: Yes, these interface functions can be merged and reformulated to eliminate duplicate functions. We can merge getBankData and getStockData into a more generic function get-stock-data, which can fetch any financial data, including bank and stock data. In addition, we can keep the prepareDataForPlotting and plotData functions, as they can also be applied to different types of financial data]
{Interface1=Interface Name:get-stock-data, ...}
{Interface2=Interface Name:prepareDataForPlotting, ...}
{Interface3=Interface Name:plotData, ...}
---- Interface Implementation -------
Using get_stock_data as examples:
def get_stock_data(name:str=’’,start_d:str=’’,end_d:str=’’,freq:str)->pd.DataFrame:
Retrieves the daily/weekly/monthly price data for a given stock code during a specific time period.
Args:
-name (str): The name of the stock.
-start_d (str): The start date in the format YYYYMMDD’.
-end_d (str): The end date in YYYYMMDD’.
-freq (str): The frequency of the price, can be daily’, weekly’, or monthly’.
Returns:
-pd.DataFrame: A dataframe that contains the daily/weekly/monthly data. The output columns contain stock_code, trade_date, open, high, low, close, pre_close, change), pct_chg,vol,amount
stock_code = get_stock_code(name)
if freq == daily’:
G.2 Cases for Real-time Request
### Request: Today is 2019-03-13. The user request is to please help me plot the stock price trend of Guizhou Maotai from January 23, 2018 to today.
### Rewritten Request: Please help me plot the closing price of Guizhou Maotai from Jan 23, 2018 to Mar 13, 2019 (Time:20180123-20190313, Location: China, Object: Guizhou Maotai stock price, Format: Line).
###Plan: task1={"stock_task": "Get the time series data of Guizhou Maotais stock price from 20180123 to 20190313"}, task2={"visualization_task": "Plot the stock price of Guizhou Maotai from 20180123 to 20190313 on a line chart"}
###Function Call:
step1={
"arg1": ["Guizhou Maotai", "20180123", "20190313", "daily"],
"function1": "get_stock_prices_data",
"output1": "result1",
"description1": "Guizhou Maotai daily stock price (open, close, high, low, etc.) time series data"},
step2={
"arg1": ["result1","close"],
"function1": "calculate_stock_index",
"output1": "result2v,
"description1": "Guizhou Maotais closing price time series data"},
step3={
"arg": ["result2", null, "line", "Guizhou Maotai closing price chart from January 23, 2018 to March 13, 2019"], "function": "plot_stock_data",
"output": "result1",
"description": "Guizhou Maotai Stock Price Line Chart"
Figure G2: Data-Copilot deploys workflows to solve users’ prediction request. It invokes three interfaces step by step and generates arguments for each interface.
Figure G3: We have listed some of Data-Copilot’s self-design interfaces, including five categories. Most of the interfaces are a combination of several simple interfaces. For example, the text marked in blue indicates that the interface includes multiple functions.
Figure G4: For complex requests about stock financial data, Data-Copilot deploys a loop workflow to solve user requests and finally outputs images and tables in parallel.
Figure G5: Data-Copilot acquires the P/E data of two stocks in parallel and plots them in a graph.
Figure G6: Data-Copilot also has access to fund data and can query the returns of all funds managed by the fund manager.
Figure G7: Data-Copilot can plot multiple indicators in a single graph by deploying workflows in parallel.
Figure G8: Data-Copilot can provide the latest financial news and company information by deploying the corresponding workflows.
Figure G9: The user interface of our system. The green box (A) is the user input panel, and the purple (B) and red parts (C) are the results returned by the system.